OpenStudy (math2400):

Here's a little math puzzle lol (: Can anyone get it? MINUTE +MONTH -------------- SEASON U must assign a number to every letter. It can be a number 0-9 To start u off, the N will be 7 So it looks like this _ _ 7 _ _ _ +_ _ _ _ _ n And another hint for the heck of it, the O is an even number

OpenStudy (anonymous):

o is 3

OpenStudy (anonymous):

She just said O is an even number 3 is an odd lol

OpenStudy (anonymous):

lol i forgot pretty tired

OpenStudy (anonymous):

so o is 4

OpenStudy (amistre64):

its just a pain to work thru ... doable, but a pain

OpenStudy (amistre64):

M I N U T E M O N T H S E A S O N M I 7 U T E M 2n 7 T H S E A S 2n 7 E+H = x+7 , x is 10 or 0 T+T+y = 2n ; y is 1 or 0 etc ...

OpenStudy (amistre64):

E+H = x+7 07 <- not useable 16 25 34 43 52 61 70 <- not useable 98 T+T+y = 2n 2T+y = 2n y = 2(n-T), y is an even number, and therefore cant be 1, 98 is unuseable T is an even number

OpenStudy (doc.brown):

\[MINUTE+MONTH=SEASON\]\[\frac{ MINUTE }{ SEASON } + \frac{ MONTH }{ SEASON } = \frac{ SEASON }{ SEASON }\]\[\frac{ MI\cancel NUT\cancel E }{ S\cancel EASO\cancel N } + \frac{ M\cancel O\cancel NTH }{ SEAS\cancel O\cancel N } = 1\]\[\frac{ MIUT }{ SASO }+\frac{ MTH }{ SEAS }=1\]\[\frac{ MIUT }{ \cancel S\cancel A\cancel SO }+\frac{ MTH }{ \cancel SE\cancel A\cancel S }=1SAS\]\[\frac{ MIUT }{ O }+\frac{ MTH }{ E } = AS^2\]

OpenStudy (anonymous):

yay